OpenStudy (anonymous):

how many values of theta satisfy the quation 3sin^2 + sin theta-2=0

OpenStudy (anonymous):

\(3sin^2(\theta)+sin(\theta)-2=0\) Let x = sin(Θ). Solve \(3x^2+x-2=0.\) But really, with no restriction on domain, there are infinite solutions.

OpenStudy (raden):

3x^2+x−2=0 (3x-2)(x+1) = 0 for zeros, x=2/3 or x=-1 sin(Θ) = 2/3 ----> Θ = arc(sin(2/3) sin(Θ) = -1 ----> Θ = 270 degrees or 3pi/2